When a car is moving 25 miles/hour, how many feet does it travel in 12 seconds?
Answer by jim_thompson5910(35256)

12 seconds = 12/(60*60) = 12/3600 = 0.00333333333333 hours


(25 mph)*(0.00333333333333 hr) = 0.08333333333 miles


So it moves roughly 0.08333333333 miles in 12 seconds


(0.08333333333 miles)*(5280 ft/1 mile) = 439.9999999824 feet


Round this to get 440 feet


So it moves roughly 440 feet in 12 seconds